# Combinatorics of Essential Sets for Positroids

### v2

@@ 3.3 Properties of essential sets
Let $\mathcal{E}^{r}$ be the ranked essential family of a positroid $\mathcal{P}$ of rank $k$ on the ground set $[n]$. Then, the following holds.
@@
    $(k,[1,n])\in\mathcal{E}^{r}$ for a unique $0<k<n$ and for each $(r,[i,j])\in\mathcal{E}^{r}$, we have
@@
    For every district pair $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{r}\setminus\{(k,[1,n])\}$ such that $[i_{1},j_{1}]\subsetneq[i_{2},j_{2}]$, we have
@@
    Let $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{r}\setminus\{(k,[1,n])\}$ such that $i_{2}<_{i_{1}}j_{2}$ and $j_{1}<_{i_{1}}j_{2}$.
@@
By construction of the essential family we have that $(k,[1,n])\in\mathcal{E}^{r}$. Moreover, if $(r,[i,j])\in\mathcal{E}^{r}\setminus\{(k,[1,n])\}$, then $10\leq r=\operatorname{rk}([i,j])=\operatorname{rk}([i+1,j])\leq|[i,j]|-$ 1.
Let $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{r}$ be distinct ranked essential sets such that $[i_{1},j_{1}]\subsetneq[i_{2},j_{2}]$. Suppose $i_{1}-1\in[i_{2},j_{2}]$ (similar argument works if $j_{1}+1\in[i_{2},j_{2}]$), then $r_{2}=\operatorname{rk}([i_{2},j_{2}])\geq\operatorname{rk}([i_{1}-1,j_{1}])=\operatorname{rk}([i_{1},j_{1}])+1>r_{1}$, hence $r_{2}-r_{1}>0$. Moreover, the following relation holds.
@@
Suppose $[i_{2},j_{2}]\neq[1,n]$ is an essential set, then the dot in the antidiagonal through $(i_{2},j_{2}-i_{2}+1)$ will be in row $h$ for $i_{2}<_{i_{2}}h<_{i_{2}}j_{2}$. If $h\in[i_{1},j_{1}]$, then the number of dots in $P_{(i_{2},j_{2}-i_{2}+1)}\setminus P_{(i_{1},j_{1}-i_{1}+1)}$ is strictly larger then $r_{2}-r_{1}$. If $h\in[i_{2},j_{2}]\setminus[i_{1},j_{1}]$, then the number of dots in $P_{(i_{2},j_{2}-i_{2}+1)}\setminus P_{(i_{1},j_{1}-i_{1}+1)}$ is strictly smaller then $|[i_{2},j_{2}]\setminus[i_{1},j_{1}]|$. Hence, one of the two inequalities from eq. 2 is strict, i.e., $r_{2}-r_{1}<|[i_{2},j_{2}]\setminus[i_{1},j_{1}]|$.
@@
Claim 1. There exists an essential set $(r,[a,b])$ containing $[i_{1},j_{1}]\cup[i_{2},j_{2}]$ of rank $\operatorname{rk}([i_{1},j_{2}])$. Indeed, if $[i_{1},j_{2}]=[1,n]$, then it is the essential set $(k,[1,n])$. Otherwise $\operatorname{rk}([i_{1},j_{2}])=\operatorname{rk}([i_{1}+1,j_{2}])=\operatorname{rk}([i_{1},j_{2}-1])$ is either equal to $k$, hence $[a,b]=[1,n]$, or is strictly smaller then $k$ and so $[a,b]$, the maximal set containing $[i_{1},j_{2}]$ with the same rank, will be an essential set.
@@ 3.4 Minimal rank conditions
As seen in the definition of connected essential sets, the rank conditions within the ranked essential family of a positroid are not minimal.
To restrict the ranked essential family to minimal conditions, we introduce a constant associated with each essential set. These constants intuitively count the number of additional dependent points in each essential set. We define them inductively as follows.
@@
Let $\mathcal{E}^{r}$ be a ranked essential family. For each ranked essential set $(r,I)\in\mathcal{E}^{r}$, let
@@
Let $\mathcal{E}^{r}$ be a ranked essential family. The core $\mathcal{C}^{r}$ of $\mathcal{E}^{r}$ is the subset given by essential sets with positive excess:
$\mathcal{C}^{r}=\{(r,I)\mid(r,I)\in\mathcal{E}^{r}\text{ and }e_{I}>0\}.$
@@
Let $\mathcal{E}^{r}$ be a ranked essential family. If an essential set $(r,[i,j])\in\mathcal{E}^{r}$ is disconnected, i.e., there exists pairwise disjoint essential sets $(r_{1},[i_{1},j_{1}]),\ldots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{r}$ such that $r=r_{1}+\dots+r_{m}+|[i,j]\setminus\bigcup_{a\in[m]}[i_{a},j_{a}]|$ and $[i,j]\supseteq\bigcup_{a\in[m]}[i_{a},j_{a}]$, then $(r,[i,j])$ is not in the core. Indeed
@@
However, in general the core $\mathcal{C}^{r}$ is strictly smaller than the family of ranked connected essential sets $\mathcal{C}\mathcal{E}^{r}$. Consider for example the positroid with ranked essential family
$\mathcal{E}^{r}=\{(2,[1,3]),(2,[3,5]),(3,[1,5]),(4,[1,6])\}.$
Then every essential set is connected but the core is $\mathcal{C}^{r}=\{(2,[1,3]),(2,[3,5]),(4,[1,6])\}$.
@@ 3.5 The retrieval algorithm
Description of Algorithm 1. The algorithm takes as input $E^{r}$, a subset of the array with non-negative integer labels, containing the required rank conditions. Its objective is to identify, if feasible, such a positroid satisfying such rank conditions. In the process, $\mathcal{S}$ represents the shaded squares and $D$ contains the position of the dots added to the array at any point during the algorithm. Additionally, the algorithm initializes $r$ to 00.
Consider a labeled square $(r,(i,j))\in E^{r}$. Define $a_{D}(i,j)=j-d_{D}(i,j)-r$. Let $h\in[i,j]$ be the smallest row, according to the order $<_{i}$, not containing a dot. We consider two cases:
@@
Once we have applied the first part to each labeled square, we proceed to fill the remaining rows. Let $k$ be the maximum label, which is usually appearing as the label of $(1,n)$. For each row $h$ not containing a dot, add a dot in the minimal square with rank $k$, i.e., the minimal $\ell$ such that $\ell$ minus the number of dots contained in $T_{(h,\ell)}$ equals $k$. The algorithm is formalized in Algorithm 1.
Input: $E^{r}$ a set of pair of integer number and labels for a square in an $n\times(n+1)$ array.
@@
$D\leftarrow\emptyset,\hskip 5.0pt\mathcal{S}\leftarrow\emptyset,\hskip 5.0ptr\leftarrow 0,\hskip 5.0ptk\leftarrow\max\{r\mid(r,(i,j))\in E^{r}\}$;
$E^{r}\leftarrow\textsf{Sorted}(E^{r})$ ;
@@
Consider the set of labeled squares $E^{r}=\{(1,(3,2)),(3,(1,5))\}$. The steps of the algorithm applied to $E^{r}$ can be seen in Figure 3.
@@
Let $E^{r}$ be a set of labeled squares from an $n\times(n+1)$ array and suppose we have a labeling of the square $(1,n)$. Then Algorithm 1 terminates and outputs a proper dotting if and only if there exists a positroid $\mathcal{P}$ satisfying the rank conditions corresponding to the input, i.e., for each $(r,(i,j))\in E^{r}$, $\operatorname{rk}([i,j+i-1])=r$ by $\mathcal{P}$ and such that the core of $\mathcal{P}$ is contained in the rank information encoded in $E^{r}$.
@@
Let $k$ denote the label of $(1,n)$. Note that since $k$ denotes the rank of the positroid, if $k$ is not the maximal label, then $E^{r}$ is not valid.
@@
($\Leftarrow$) Suppose the rank conditions in $E^{r}$ can be satisfied by a positroid as in the statement. We want to prove that the output of the algorithm is a proper dotting. Let $\mathcal{D}$ be a maximal proper dotting such that $r_{\mathcal{D}}(i,j)=r$ for every $(r,(i,j))\in E^{r}$ and the core of the bounded affine permutation corresponding to $\mathcal{D}$ is contained in the rank information given by $E^{r}$. We want to prove that at every step of the algorithm, the dotting $D$ produced is a subset of $\mathcal{D}$.
@@
In the first loop, for each $(r,(i,j))\in E^{r}$ considered in increasing rank order, we add dots inside the triangle $T_{(i,j)}$, to restrict the rank of $[i,j+i-1]$ to be at maximum $r$. Let $D$ be the dotting at this step and suppose $D$ is a subset of $\mathcal{D}$. We enter the loop if $j-d_{D}(i,j)>r$, that is whenever the number of dots in $T_{(i,j)}$ is not high enough. Since the quantity $j-r-d_{D}(i,j)$ is the excess of $[i,j+i-1]$, this is equivalent to the requirement that $(i,j)$ corresponds to a cyclic interval in the core of $\mathcal{D}$. Let $h$ be the first empty row starting from $i$. Since $D$ can be extended to $\mathcal{D}$ and $r_{\mathcal{D}}(i,j)=r$, then $h\in[i,j+i-1]$. Let $\ell$ be such that $(h,\ell)\in\mathcal{D}$. Note that $r_{\mathcal{D}}(h,j+i-h)=r$. Indeed the dots that have been added so far $D$ in rows $i$ to $i+h-2$ lie inside $T_{(i,j)}$. Suppose not, let the first dot in rows $i$ to $i+h-2$ lying outside of $T_{(i,j)}$ be in square $(h_{2},\ell_{2})$, added while considering some square in $E^{r}$ labeled by $r_{2}\leq r$. Then $i+j-h-d_{D}(h_{2},i+j-h)<r_{2}$, hence $d_{D}(h_{2},i+j-h)>i+j-h-r_{2}\geq i+j-h-r$. Since all the rows between $i$ and $h_{2}$ have been filled by assumption with dots inside $T_{(i,j)}$, this would imply that $d_{D}(i,j)>h-i+i+j-h-r=j-r$, against the assumption that $a=j-r-d_{D}(i,j)>0$. It follows that $r_{\mathcal{D}}(h,\ell)=r$ and $\ell\in T_{(i,j)}$. Moreover, $\ell^{\prime}=\min\{\beta\mid\beta-1-d_{D}(h,\beta)=r\}$ is such that $\ell=\ell^{\prime}$ by rank maximality of $\mathcal{D}$. Hence the algorithm does not produce an error and $D\cup\{(h,\ell^{\prime})\}\subseteq\mathcal{D}$.
@@
In the second for-loop, we add the dots outside of each $T_{(i,j)}$ considered in the first loop. Suppose $D$ is the dotting constructed so far and it is a subset of $\mathcal{D}$ and let $h$ be the row we are considering. Let $\ell^{\prime}=\min\{\beta\mid\beta-1-d_{D}(h,\beta)=k\}$, where $k$ is the label of the square $(1,n)$. Then, since for every $(r,(i,j))\in E^{r}$, $d_{D}(i,j)\geq j-r$, $\ell^{\prime}$ will lie outside of every triangle $T_{(i,j)}$. This guarantees that the rank of the square $(i,j)$ with respect to the dotting we are constructing is actually equal to $r$ for each $(r,(i,j))\in E^{r}$. Let $\ell$ be such that $(h,\ell)\in\mathcal{D}$. By rank maximality of $\mathcal{D}$, $r_{\mathcal{D}}(h,\ell)=k$ and $\ell\geq\ell^{\prime}$. Moreover, suppose by contradiction that $\ell>\ell^{\prime}$. Then since $r_{\mathcal{D}}(h,\ell)=r_{\mathcal{D}}(h,\ell^{\prime})=k$, $\mathcal{D}$ contains a dot in the antidiagonal containing $(h,\ell^{\prime})$ in the rows between $h$ and $h+\ell^{\prime}-1$. This is not possible by definition of $\ell^{\prime}$. Hence $\ell=\ell^{\prime}$, the dotting produced is still a subset of $\mathcal{D}$ and the algorithm does not produce any error at this step.
@@
Since we assume $\mathcal{D}$ to be a maximal proper dotting satisfying the rank conditions in $E^{r}$ and we proved that the dotting $D$ constructed by the algorithm is equal to $\mathcal{D}$, the last two checks will be satisfied by $D$, hence the algorithm will terminate and output $D$.
@@
($\Rightarrow$) Suppose now that the algorithm terminates and outputs a dotting, denoted as $D$. Note that $D$ must satisfy the rank conditions specified by $E^{r}$, as the last step in the algorithm checks for this. Moreover, at every step, we choose a row $h$ that does not already contain a dot. As a consequence, at the end of the algorithm, all the boxes will be shaded if and only if no two dots lie in the same antidiagonal. Hence, $D$ is a proper dotting satisfying the rank conditions given in the input; that is, there exists a positroid satisfying rk ([$\textrm{rk }([i,j+i-1])=r$ for every $(r,(i,j))\in E^{r}$. ∎

### v3

@@ 3.3 Properties of essential sets
Let $\mathcal{E}^{\mathrm{r}}$ be the ranked essential family of a positroid $\mathcal{P}$ of rank $k$ on the ground set $[n]$. Then, the following holds.
@@
    $(k,[1,n])\in\mathcal{E}^{\mathrm{r}}$ and for each $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}$, we have
@@
    For every distinct pair $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{\mathrm{r}}\setminus\{(k,[1,n])\}$ such that $[i_{1},j_{1}]\subsetneq[i_{2},j_{2}]$, we have
@@
    Let $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{\mathrm{r}}\setminus\{(k,[1,n])\}$ such that $i_{2}<_{i_{1}}j_{2}$ and $j_{1}<_{i_{1}}j_{2}$.
@@
By construction of the essential family we have that $(k,[1,n])\in\mathcal{E}^{\mathrm{r}}$. Moreover, if $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}\setminus\{(k,[1,n])\}$, then $10\leq r=\operatorname{rk}([i,j])=\operatorname{rk}([i+1,j])\leq|[i,j]|-$ 1.
Let $(r_{1},[i_{1},j_{1}]),(r_{2},[i_{2},j_{2}])\in\mathcal{E}^{\mathrm{r}}$ be distinct ranked essential sets such that $[i_{1},j_{1}]\subsetneq[i_{2},j_{2}]$. Suppose $i_{1}-1\in[i_{2},j_{2}]$ (similar argument works if $j_{1}+1\in[i_{2},j_{2}]$), then $r_{2}=\operatorname{rk}([i_{2},j_{2}])\geq\operatorname{rk}([i_{1}-1,j_{1}])=\operatorname{rk}([i_{1},j_{1}])+1>r_{1}$, hence $r_{2}-r_{1}>0$. Moreover, the following relation holds.
@@
Suppose $[i_{2},j_{2}]\neq[1,n]$ is an essential set, then the dot in the antidiagonal through $(i_{2},j_{2}-i_{2}+1)$ will be in row $h$ for $i_{2}<_{i_{2}}h<_{i_{2}}j_{2}$. If $h\in[i_{1},j_{1}]$, then the number of dots in $P_{(i_{2},j_{2}-i_{2}+1)}\setminus P_{(i_{1},j_{1}-i_{1}+1)}$ is strictly larger than $r_{2}-r_{1}$. If $h\in[i_{2},j_{2}]\setminus[i_{1},j_{1}]$, then the number of dots in $P_{(i_{2},j_{2}-i_{2}+1)}\setminus P_{(i_{1},j_{1}-i_{1}+1)}$ is strictly smaller than $|[i_{2},j_{2}]\setminus[i_{1},j_{1}]|$. Hence, one of the two inequalities from eq. 2 is strict, i.e. $r_{2}-r_{1}<|[i_{2},j_{2}]\setminus[i_{1},j_{1}]|$.
@@
Claim 1. There exists an essential set $(r,[a,b])$ containing $[i_{1},j_{1}]\cup[i_{2},j_{2}]$ of rank $\operatorname{rk}([i_{1},j_{2}])$. Indeed, if $[i_{1},j_{2}]=[1,n]$, then it is the essential set $(k,[1,n])$. Otherwise $\operatorname{rk}([i_{1},j_{2}])=\operatorname{rk}([i_{1}+1,j_{2}])=\operatorname{rk}([i_{1},j_{2}-1])$ is either equal to $k$, hence $[a,b]=[1,n]$, or is strictly smaller than $k$ and so $[a,b]$, the maximal set containing $[i_{1},j_{2}]$ with the same rank, will be an essential set.
@@ 3.4 Minimal rank conditions
As seen in the definition of connected essential sets, the rank conditions within the ranked essential family of a positroid are not minimal. To restrict the ranked essential family to minimal conditions, we introduce a constant associated with each essential set. These constants intuitively count the number of additional dependent points in each essential set. We define them inductively as follows.
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a ranked essential family. For each ranked essential set $(r,I)\in\mathcal{E}^{\mathrm{r}}$, let
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a ranked essential family. The core $\mathcal{C}^{\mathrm{r}}$ of $\mathcal{E}^{\mathrm{r}}$ is the subset given by essential sets with positive excess:
$\mathcal{C}^{\mathrm{r}}=\{(r,I)\mid(r,I)\in\mathcal{E}^{\mathrm{r}}\text{ and}e_{I}>0\}.$
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a ranked essential family. If an essential set $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}$ is disconnected, i.e. there exists pairwise disjoint essential sets $(r_{1},[i_{1},j_{1}]),\ldots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{\mathrm{r}}$ such that $r=r_{1}+\dots+r_{m}+|[i,j]\setminus\bigcup_{a\in[m]}[i_{a},j_{a}]|$ and $[i,j]\supseteq\bigcup_{a\in[m]}[i_{a},j_{a}]$, then $(r,[i,j])$ is not in the core. Indeed
@@
However, in general the core $\mathcal{C}^{\mathrm{r}}$ is strictly smaller than the family of ranked connected essential sets $\mathcal{C}\mathcal{E}^{\mathrm{r}}$. Consider for example the positroid with ranked essential family
$\mathcal{E}^{\mathrm{r}}=\{(2,[1,3]),(2,[3,5]),(3,[1,5]),(4,[1,6])\}.$
Then every essential set is connected but the core is $\mathcal{C}^{\mathrm{r}}=\{(2,[1,3]),(2,[3,5]),(4,[1,6])\}$.
@@ 3.5 The retrieval algorithm
Description of Algorithm 1. The algorithm takes as input $E^{\mathrm{r}}$, a subset of the array with non-negative integer labels, containing the required rank conditions. Its objective is to identify, if feasible, a positroid satisfying such rank conditions. In the process, $\mathcal{S}$ represents the shaded squares and $D$ contains the position of the dots added to the array at any point during the algorithm. Additionally, the algorithm initializes $r$ to 00.
Consider a labeled square $(r,(i,j))\in E^{\mathrm{r}}$. Define $a_{D}(i,j)=j-d_{D}(i,j)-r$. Let $h\in[i,j]$ be the smallest row, according to the order $<_{i}$, not containing a dot. We consider two cases:
@@
Once we have applied the first part to each labeled square, we proceed to fill the remaining rows. Let $k$ be the maximum label, which is usually appearing as the label of $(1,n)$. For each row $h$ not containing a dot, add a dot in the minimal square with rank $k$, i.e. the minimal $\ell$ such that $\ell$ minus the number of dots contained in $T_{(h,\ell)}$ equals $k$. The algorithm is formalized in Algorithm 1.
Input: $E^{\mathrm{r}}$ a set of pair of integer number and labels for a square in an $n\times(n+1)$ array.
@@
$D\leftarrow\emptyset,\hskip 5.0pt\mathcal{S}\leftarrow\emptyset,\hskip 5.0ptr\leftarrow 0,\hskip 5.0ptk\leftarrow\max\{r\mid(r,(i,j))\in E^{\mathrm{r}}\}$;
$E^{\mathrm{r}}\leftarrow\textsf{Sorted}(E^{\mathrm{r}})$ ;
@@
Consider the set of labeled squares $E^{\mathrm{r}}=\{(1,(3,2)),(3,(1,5))\}$. The steps of the algorithm applied to $E^{\mathrm{r}}$ can be seen in Figure 3.
@@
Let $E^{\mathrm{r}}$ be a set of labeled squares from an $n\times(n+1)$ array and suppose we have a labeling of the square $(1,n)$. Then Algorithm 1 terminates and outputs a proper dotting if and only if there exists a positroid $\mathcal{P}$ satisfying the rank conditions corresponding to the input, i.e. for each $(r,(i,j))\in E^{\mathrm{r}}$, $\operatorname{rk}([i,j+i-1])=r$ by $\mathcal{P}$ and such that the core of $\mathcal{P}$ is contained in the rank information encoded in $E^{\mathrm{r}}$.
@@
Let $k$ denote the label of $(1,n)$. Note that since $k$ denotes the rank of the positroid, if $k$ is not the maximal label, then $E^{\mathrm{r}}$ is not valid.
@@
($\Leftarrow$) Suppose the rank conditions in $E^{\mathrm{r}}$ can be satisfied by a positroid as in the statement. We want to prove that the output of the algorithm is a proper dotting. Let $\mathcal{D}$ be a maximal proper dotting such that $r_{\mathcal{D}}(i,j)=r$ for every $(r,(i,j))\in E^{\mathrm{r}}$ and the core of the bounded affine permutation corresponding to $\mathcal{D}$ is contained in the rank information given by $E^{\mathrm{r}}$. We want to prove that at every step of the algorithm, the dotting $D$ produced is a subset of $\mathcal{D}$.
@@
In the first loop, for each $(r,(i,j))\in E^{\mathrm{r}}$ considered in increasing rank order, we add dots inside the triangle $T_{(i,j)}$, to restrict the rank of $[i,j+i-1]$ to be at maximum $r$. Let $D$ be the dotting at this step and suppose $D$ is a subset of $\mathcal{D}$. We enter the loop if $j-d_{D}(i,j)>r$, that is whenever the number of dots in $T_{(i,j)}$ is not high enough. Since the quantity $j-r-d_{D}(i,j)$ is the excess of $[i,j+i-1]$, this is equivalent to the requirement that $(i,j)$ corresponds to a cyclic interval in the core of $\mathcal{D}$. Let $h$ be the first empty row starting from $i$. Since $D$ can be extended to $\mathcal{D}$ and $r_{\mathcal{D}}(i,j)=r$, then $h\in[i,j+i-1]$. Let $\ell$ be such that $(h,\ell)\in\mathcal{D}$. Note that $r_{\mathcal{D}}(h,j+i-h)=r$. Indeed the dots that have been added so far $D$ in rows $i$ to $i+h-2$ lie inside $T_{(i,j)}$. Suppose not, let the first dot in rows $i$ to $i+h-2$ lying outside of $T_{(i,j)}$ be in square $(h_{2},\ell_{2})$, added while considering some square in $E^{\mathrm{r}}$ labeled by $r_{2}\leq r$. Then $i+j-h-d_{D}(h_{2},i+j-h)<r_{2}$, hence $d_{D}(h_{2},i+j-h)>i+j-h-r_{2}\geq i+j-h-r$. Since all the rows between $i$ and $h_{2}$ have been filled by assumption with dots inside $T_{(i,j)}$, this would imply that $d_{D}(i,j)>h-i+i+j-h-r=j-r$, against the assumption that $a=j-r-d_{D}(i,j)>0$. It follows that $r_{\mathcal{D}}(h,\ell)=r$ and $\ell\in T_{(i,j)}$. Moreover, $\ell^{\prime}=\min\{\beta\mid\beta-1-d_{D}(h,\beta)=r\}$ is such that $\ell=\ell^{\prime}$ by rank maximality of $\mathcal{D}$. Hence the algorithm does not produce an error and $D\cup\{(h,\ell^{\prime})\}\subseteq\mathcal{D}$.
@@
In the second for-loop, we add the dots outside of each $T_{(i,j)}$ considered in the first loop. Suppose $D$ is the dotting constructed so far and it is a subset of $\mathcal{D}$ and let $h$ be the row we are considering. Let $\ell^{\prime}=\min\{\beta\mid\beta-1-d_{D}(h,\beta)=k\}$, where $k$ is the label of the square $(1,n)$. Then, since for every $(r,(i,j))\in E^{\mathrm{r}}$, $d_{D}(i,j)\geq j-r$, $\ell^{\prime}$ will lie outside of every triangle $T_{(i,j)}$. This guarantees that the rank of the square $(i,j)$ with respect to the dotting we are constructing is actually equal to $r$ for each $(r,(i,j))\in E^{\mathrm{r}}$. Let $\ell$ be such that $(h,\ell)\in\mathcal{D}$. By rank maximality of $\mathcal{D}$, $r_{\mathcal{D}}(h,\ell)=k$ and $\ell\geq\ell^{\prime}$. Moreover, suppose for contradiction that $\ell>\ell^{\prime}$. Then since $r_{\mathcal{D}}(h,\ell)=r_{\mathcal{D}}(h,\ell^{\prime})=k$, $\mathcal{D}$ contains a dot in the antidiagonal containing $(h,\ell^{\prime})$ in the rows between $h$ and $h+\ell^{\prime}-1$. This is not possible by definition of $\ell^{\prime}$. Hence $\ell=\ell^{\prime}$, the dotting produced is still a subset of $\mathcal{D}$ and the algorithm does not produce any error at this step.
@@
Since we assume $\mathcal{D}$ to be a maximal proper dotting satisfying the rank conditions in $E^{\mathrm{r}}$ and we proved that the dotting $D$ constructed by the algorithm is equal to $\mathcal{D}$, the last two checks will be satisfied by $D$, hence the algorithm will terminate and output $D$.
@@
($\Rightarrow$) Suppose now that the algorithm terminates and outputs a dotting, denoted as $D$. Note that $D$ must satisfy the rank conditions specified by $E^{\mathrm{r}}$, as the last step in the algorithm checks for this. Moreover, at every step, we choose a row $h$ that does not already contain a dot. As a consequence, at the end of the algorithm, all the boxes will be shaded if and only if no two dots lie in the same antidiagonal. Hence, $D$ is a proper dotting satisfying the rank conditions given in the input; that is, there exists a positroid satisfying rk ([$\textrm{rk }([i,j+i-1])=r$ for every $(r,(i,j))\in E^{\mathrm{r}}$. ∎